OpenStudy (anonymous):

Find the missing coordinates in the following solution for the given equation: Y = X2 + 2X + 1 : (-3,?) (X2 is supposed to be X squared but I haven't figured out how to do that on the puter)

OpenStudy (anonymous):

simplify by putting x=-3 hence we have y=(-3)^2 +2(-3)+1=9-6+1=4 hence the missing cordinate is 4

OpenStudy (anonymous):

Good Lord.. was that seriously all there was to that? After all the other stuff I have been working on tonight, I seriously was making that out to be a whole lot more complicated then it actually was.

OpenStudy (anonymous):

Here is a really stupid question... again, its late and I have been working tirelessly on this for more then 15 hours, so I know I know the answer, but I am drawing a blank... -6 (2/3)... is that = to: -12/3

OpenStudy (anonymous):

yup -6 (2/3)... is that = to: -12/3=-4

OpenStudy (anonymous):

Here is the actual question I am trying to answer: Y = -6x - 5 : (2/3, ?) Y= -6(2/3) - 5 12/3 - 15/3 Y = -3

OpenStudy (anonymous):

Y= -6(2/3) - 5=-(12/3) -5 =-4 -5 =-9

OpenStudy (anonymous):

Damn.. I was way off on that one

OpenStudy (anonymous):

It is the multiplication of the fractions... it gets me every time